OpenStudy (anonymous):

let f(x)=-2x-7 and g(x)=-4x+3 find (f o g ) (-5) help please :)

OpenStudy (anonymous):

sorry its (f o g ) (x)

OpenStudy (amistre64):

same concept :) we just need to insert g(x) INTO f(x) f(x) = -2x - 7 f(g(x)) = -2 (g(x)) - 7 = -2(-4x+3) - 7

OpenStudy (anonymous):

8x-1?

OpenStudy (amistre64):

8x -6 -7 = 8x -14

OpenStudy (amistre64):

gotta keep your eye on those signs :)

OpenStudy (anonymous):

8x-13? thats one of my answer choices..8x-14 isn't

OpenStudy (amistre64):

doh!!, i forgot how to add lol

OpenStudy (anonymous):

lOl!!!

OpenStudy (amistre64):

8x - 13 will be fine then :)